A random sample from a company’s very extensive files show that orders for a certain piece of machinery were filled, respectively in 12, 10, 17, 14, 13, 18, 11 and 9 days. Use the level of significance 0.05 to test the claim that on the average such orders are filled in 9.5 days against the alternative hypothesis µ>9.5.
Answer #1

H0: = 9.5

Ha: > 9.5

From Sample data,

= 13 . S = 3.2071

Test Statistic :-
t = ( X̅ - µ ) / (S / √(n) )
t = ( 13 - 9.5 ) / ( 3.2071 / √(8) )
t = 3.09

Test Criteria :-
Reject null hypothesis if t > t(α, n-1)
Critical value t(α, n-1) = t(0.05 , 8-1) = 1.895

t > t(α, n-1) = 3.09 > 1.895
Result :- Reject null hypothesis